An experiment calls for a 5% salt solution. Only a 3% salt solution and an 6% salt solution are available. How many gallons of each must be combined to obtain 30 gallons of a 5% salt solution?

Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

0.03x+0.06(30-x) = 0.05(30)
0.03x+1.8-0.06x = 1.5
-0.03x = -0.3 Divide, and remember that -/- = +
x = 10
You'll need 10 gallon of the 3% solution and 90-10 = 80 gallon of the 6% solution to obtain 90 gallon of a 5% solution

Answer by ikleyn(52834)

.
An experiment calls for a 5% salt solution. Only a 3% salt solution and an 6% salt solution are available.
How many gallons of each must be combined to obtain 30 gallons of a 5% salt solution?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

0.03x+0.06(30-x) = 0.05(30)
0.03x+1.8-0.06x = 1.5
-0.03x = -0.3 Divide.
x = 10
You'll need 10 gallon of the 3% solution and 30-10 = 20 gallon of the 6% solution to obtain 30 gallon of a 5% solution